# HiLo: A Learning Framework for Generalized Category Discovery Robust to Domain Shifts

### v1

@@ Abstract
Generalized Category Discovery (GCD) is a challenging task in which, given a partially labelled dataset, models must categorize all unlabelled instances, regardless of whether they come from labelled categories or from new ones. In this paper, we challenge a remaining assumption in this task: that all images share the same domain. Specifically, we introduce a new task and method to handle GCD when the unlabelled data also contains images from different domains to the labelled set. Our proposed ‘HiLo’ networks extract High-level semantic and Low-level domain features, before minimizing the mutual information between the representations. Our intuition is that the clusterings based on domain information and semantic information should be independent. We further extend our method with a specialized domain augmentation tailored for the GCD task, as well as a curriculum learning approach. Finally, we construct a benchmark from corrupted fine-grained datasets as well as a large-scale evaluation on DomainNet with real-world domain shifts, reimplementing a number of GCD baselines in this setting. We demonstrate that HiLo outperforms SoTA category discovery models by a large margin on all evaluations.
@@ 1 Introduction
The task of category discovery [1] has recently gained substantial interest in the computer vision community [2, 3, 4, 5, 6, 7]. The task is to leverage knowledge from a number of labelled images, in order to discover and cluster images from novel classes in unlabelled data. Such a task naturally occurs in many practical settings; from products in a supermarket, to animals in the wild, to street objects for an autonomous vehicle. Specifically, Generalized Category Discovery (GCD) [8] has recently emerged as a challenging variant of the problem in which the unlabelled data can contain both instances from ‘seen’ and ‘unseen’ classes. As such, the problem is succinctly phrased as: “given a dataset, some of which is labelled, categorise all unlabelled instances (whether or not they come from labelled classes)”. This is more realistic and challenging than conventional Semi-Supervised Learning (SSL) [9], which aims to learn a strong model to recognize objects using both labelled and unlabelled data, under the assumption that the unlabelled samples belong to the same classes as the labelled data.
In this paper, we challenge a key, but often ignored, assumption in this setting: GCD methods still assume that all instances in the unlabelled set come from the same domain as the labelled data. In practise, unlabelled images may not only contain novel categories, but also exhibit low-level covariate shift [10, 11]. It has long been established that the performance of image classifiers degrades substantially in the presence of such shifts [12, 13, 14] and, indeed, we find that existing GCD models perform poorly in such a setting. Compared to related literature in, for instance, domain adaptation [15, 16, 17] or domain generalization [18, 19] the task proposed here presents a dual challenge: models must be robust to the low-level covariate shift while remaining sensitive to semantic novelty.
@@
To tackle these problems, we introduce the ‘HiLo’ architecture and learning framework. The HiLo architecture extracts both ‘low-level’ (early layer) and ‘high-level’ (late layer) features from a vision transformer [20]. While extracting features at multiple stages of the network has been performed in domain adaptation [21, 22, 23], we further introduce an explicit loss term to minimise mutual information between the two sets of features (Section 3.2.1). The intuition is that the covariate and semantic information in the data is (by definition) independent, and that the inductive bias of deep architectures is likely to represent low-level covariate information in early layers, and abstract semantic information in later ones [24, 25]. Next, we take inspiration from a strong method from the domain adaptation field, PatchMix [17], which works by performing mixup augmentation in the embedding space of a pretrained transformer. While naive application of this method does not account for semantic novelty in unlabelled data, we extend the PatchMix objective to allow training with both a self-supervised contrastive objective (Section 3.2.2), and a semantic clustering loss (Section 3.2.2). With these changes, the PatchMix style augmentation is tailored to leverage both the labelled and unlabelled data available in the GCD setting. Our ‘HiLo’ feature design in our framework enables the model to disentangle domain and semantic features, while patch mixing allows the model to bridge the domain gap among images and focus more on determining the semantic shifts. Therefore, we introduce the patch mixing idea into our ‘HiLo’ framework, equipping it with a strong capability to discover novel categories from unlabelled images in the presence of domain shifts.
@@
Finally, we find that curriculum learning [26, 27, 28] is particularly applicable to the setting introduced in this work (Section 3.2.3). Specifically, the quality of the learning signal differs substantially across different partitions of the data: from a clean supervised signal on the labelled set; to unsupervised signals from unlabelled data which may or may not come from the same domain and categories. It is non-trivial to train a GCD model to discover novel categories in the presence of both domain shifts and semantic shifts in the unlabelled data. To address this challenge, we introduce a curriculum learning approach which gradually increases the sampling probability weight of samples predicted as from unknown domains, as training proceeds. Our sequential learning process prioritizes the discovery of semantic categories initially and progressively enhances the model’s ability to handle covariate shifts, which cannot be achieved by simply adopting existing domain adaptation methods.
@@
To evaluate out models, we construct the ‘SSB-C’ benchmark suite – based on the recent Semantic Shift Benchmark (SSB) [29] – with domain shifts introduced by synthetic corruptions following ImageNet-C [30]. On this benchmark, as well as on a large-scale DomainNet evaluation with real data [31], we also reimplement a range of performant baselines from the category discovery literature. We find that, on both benchmarks, our method substantially outperforms all existing category discovery models [8, 5, 1, 4].
@@
In summary, we make the following key contributions: (i) We formalize a challenging open-world task for category discovery in the presence of domain shifts; (ii) We develop a new method, HiLo, which disentangles covariate and semantic features to tackle the problem, extending state-of-the-art methods from the domain adaptation literature; (iii) We reimplement a range of category discovery models on a benchmark suite containing both fine-grained and coarse-grained datasets, with real and synthetic corruptions. (iv) We demonstrate that, on all datasets, our method substantially outperforms current state-of-the-art category discovery methods with finetuned hyperparamters.
@@ 2 Related work
Category discovery was firstly studied as novel category discovery (NCD) [1] and recently extended to generalized category discovery (GCD) [8]. GCD extends NCD by including unlabelled images from both labelled and novel categories. Many successful NCD methods have been proposed (e.g., DTC [1], RankStats [2, 3], WTA [6], DualRank [7], OpenMix [32], NCL [33], UNO [4], Spacing loss [34]), they do not address domain shifts. Recent work [35] considers domain shifts in NCD with labelled target domain images. We focus on GCD without any labelled instances from new domains, where unlabelled images may come from multiple novel domains. For GCD, [8] fine-tunes a ViT model using DINO [36] and semi-supervised $k$-means clustering. ORCA [37] enhances intra-class separability with adaptive margin loss. CiPR [38] uses hierarchical clustering and positive samples for representation learning. SimGCD [5] employs entropy regularization for improved performance. Other methods include [39, 40, 41, 42, 43]. DCCL [39] dynamically updates visual conceptions. PromptCAL [40] refines affinity graphs in vision transformers. GPC [42] uses a GMM-based method for representation learning and category estimation. $\mu$-GCD [41] applies a student-teacher mechanism. The study of category discovery is attracting increasing efforts, expanding to different challenging scenarios, e.g., the presence of long-tailed data distribution [44, 45], collaborative discovery of novel classes across distributed datasets while maintaining data privacy [46, 47], continual category discovery without forgetting the previously learned categories [48, 49, 50], etc. The challenge of domain shifts has been considered in the NCD setting by [51] under two critical assumptions, the unlabelled samples are from a disjoint set of categories than the labelled ones, and the unlabelled samples are from a single new domain w.r.t. the labelled data, rendering this setting less practical. In contrast, in this paper, we consider the problem of GCD with domain shifts, relaxing semantic constraints to the GCD setting while taking into account the mixed domains in the unlabelled data. This more realistic and challenging problem, to our knowledge, has not been explored in the literature.
@@
Unsupervised domain adaptation (UDA) adapts models from a source domain to a target domain, with labelled data from the former and unlabelled data from the latter. UDA methods are categorized into moment matching [13, 52, 53, 14] and adversarial learning [12, 54, 55] methods. DANN, FGDA, DADA are popular examples using a min-max game. MCD and SWD implicitly use adversarial learning with $L_{1}$ distance and sliced Wasserstein discrepancy, respectively. [56] extracts foreground semantics using class activation maps and mix them with randomly cropped patches from other domains. Mixstyle [25] enhances model generalizability by probabilistically mixing instance-level feature statistics of training samples from different source domains. EFDM [57] applies the Exact Histogram Matching (EHM) algorithm, Sort-Matching, to the empirical Cumulative Distribution Functions (eCDFs) of image features for cross-domain generalization. CGDM [15] leverages cross-domain gradient discrepancy, while [16] couples NWD with a single task-specific classifier with implicit K-Lipschitz constraint. PMTrans [17] aligns the source and target domains with the intermediate domain by employing semi-supervised mixup losses in both feature and label spaces. UDA typically assumes that the source and target domains have a shared label space, while [58] introduces the task of universal domain adaptation, which aims to learn a domain-invariant representation for data from different domains with partially overlapping label spaces. UniOT [59] leverages the optimal transport for partial alignment with adaptive filling to encourage global discrimination and local consistency of samples for universal domain adaptation.
@@ 3 HiLo networks for GCD with domain shifts
Problem statement. We define Generalized Category Discovery with domain shifts as the task of classifying images from mixed domains $\Omega=\Omega^{a}\cup\Omega^{b}$ (where $\Omega^{a}\cap\Omega^{b}=\emptyset$ and $\Omega^{b}$ may contain multiple domains in practise), only having access to partially labelled samples from domain $\Omega^{a}$. The goal is to assign class labels to the remaining images, whose categories and domains may be seen or unseen in the labelled images. Formally, let $\mathcal{D}$ be an open-world dataset consisting of a labelled set $\mathcal{D}^{l}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N_{l}}\subset\mathcal{X}^{l}\times\mathcal{Y}^{l}$ and an unlabelled set $\mathcal{D}^{u}=\{\boldsymbol{x}_{i}\}_{i=1}^{N_{u}}\subset\mathcal{X}^{u}$. The label space for labelled samples is $\mathcal{Y}^{l}=\mathcal{C}_{1}$ and for unlabelled samples is $\mathcal{Y}^{u}=\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$, where $\mathcal{C}$, $\mathcal{C}_{1}$, and $\mathcal{C}_{2}$ represent the label sets for ‘All’, ‘Old’, and ‘New’ categories, respectively. It is important to note that $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$. The objective of GCD is to classify all unlabelled images in $\mathcal{D}^{u}$ (from either $\Omega^{a}$ or $\Omega^{b}$) using only the labels in $\mathcal{D}^{l}$. This is different from the setting of NCD with domain shift and GCD, which assumes $\mathcal{Y}^{l}\cap\mathcal{Y}^{u}=\emptyset$ for the former and $\Omega^{a}=\Omega^{b}$ with singleton cardinalities for the latter. For notation simplicity, hereafter we omit the subscript $i$ for each image $\boldsymbol{x}_{i}$.
@@ 3.1 Background: SimGCD
SimGCD [5] is a representative end-to-end baseline for GCD, which integrates two primary losses for representation learning and parametric classification: (1) a contrastive loss $\mathcal{L}^{rep}$ based on InfoNCE [60] is applied for the representation learning of the feature backbone; and (2) a cross-entropy loss $\mathcal{L}^{cls}$ for training a cosine classification head [61], utilizing different image views as pseudo-labels for one another. Following [8], SimGCD employs the ViT model as the backbone containing $m$ Transformer layers. Let $\mathcal{F}$ be the feature extractor consisting of these $m$ layers and $\mathcal{H}$ be a projection head. For an input image $\boldsymbol{x}$, a $\ell_{2}$-normalised feature can be obtained by $\boldsymbol{z}=\mathcal{H}(\mathcal{F}(\varphi(\boldsymbol{x})))$, where $\varphi$ is a standard embedding layer before the multi-head attention layers in the ViT model. The representation loss is
$\mathcal{L}^{rep}(\boldsymbol{x})=-\frac{1}{|\mathcal{P}(\boldsymbol{x})|}\sum_{\boldsymbol{z}^{+}\in\mathcal{P}(\boldsymbol{x})}\log\sigma(\boldsymbol{z}\cdot\boldsymbol{z}^{+};\tau),$ (1)
@@
where $\sigma(\cdot;\tau)$ is the softmax operation with a temperature $\tau$ for scaling and $\mathcal{P}(\boldsymbol{x})$ denotes the positive feature set for each $\boldsymbol{x}$. Suppose we sample a batch $\mathcal{B}$, which contains labelled images and unlabelled images, denoted as $\mathcal{B}^{l}$ and $\mathcal{B}^{u}$, respectively. For each $\boldsymbol{x}\in\mathcal{B}$ (either a labelled or unlabelled image), $\mathcal{P}(\boldsymbol{x})$ contains only the feature of a different view of the same image. For each $\boldsymbol{x}\in\mathcal{B}^{l}$, an additional $\mathcal{P}(\boldsymbol{x})$ including features of other images from the same class and the feature of a different view of the same image is also used for supervised constrastive learning. Likewise, the classification loss can be written as
$\mathcal{L}^{cls}(\boldsymbol{x})=-\sum_{\boldsymbol{w}\in\boldsymbol{W}}\boldsymbol{q}\log\sigma(\hat{\boldsymbol{z}}\cdot\boldsymbol{w};\tau),$ (2)
@@
where $\boldsymbol{W}$ is a set of prototypes and each vector $\boldsymbol{w}$ in $\boldsymbol{W}$ represents a $\ell_{2}$-normalised learnable class prototype. $\hat{\boldsymbol{z}}$ is the $\ell_{2}$-normalised vector of $\mathcal{F}(\varphi(\boldsymbol{x}))$. For each $\boldsymbol{x}\in\mathcal{B}$, $\boldsymbol{q}$ is the pseudo-label from a sharpened prediction of a different view of the same image. For each $\boldsymbol{x}\in\mathcal{B}^{l}$, an additional $\boldsymbol{q}$ as the one-hot ground-truth vector is also used for supervised learning. Let $\mathcal{L}^{r,c}$ be the summation of $\mathcal{L}^{rep}$ and $\mathcal{L}^{cls}$ for simplification, the overall loss can then be written as:
$\mathcal{L}_{sim}=\lambda\sum_{\boldsymbol{x}\in\mathcal{B}}\mathcal{L}^{r,c}(\boldsymbol{x})+(1-\lambda)\sum_{\boldsymbol{x}\in\mathcal{B}^{l}}\mathcal{L}^{r,c}(\boldsymbol{x})+\epsilon\Delta,$ (3)
@@
where $\mathcal{B}^{l}$ denotes the subset of labelled samples in the current mini-batch, and $\Delta$ is an entropy maximization term to prevent pseudo-label collapse [62]. Finally, $\lambda$ and $\epsilon$ are hyperparameters, and we refer to the original work for further details [5].

### v2

@@ Abstract
Generalized Category Discovery (GCD) is a challenging task in which, given a partially labelled dataset, models must categorize all unlabelled instances, regardless of whether they come from labelled categories or from new ones. In this paper, we challenge a remaining assumption in this task: that all images share the same domain. Specifically, we introduce a new task and method to handle GCD when the unlabelled data also contains images from different domains to the labelled set. Our proposed ‘HiLo’ networks extract High-level semantic and Low-level domain features, before minimizing the mutual information between the representations. Our intuition is that the clusterings based on domain information and semantic information should be independent. We further extend our method with a specialized domain augmentation tailored for the GCD task, as well as a curriculum learning approach. Finally, we construct a benchmark from corrupted fine-grained datasets as well as a large-scale evaluation on DomainNet with real-world domain shifts, reimplementing a number of GCD baselines in this setting. We demonstrate that HiLo outperforms SoTA category discovery models by a large margin on all evaluations. Project page: https://visual-ai.github.io/hilo/
@@ 1 Introduction
Category discovery Han et al. (2019) has recently gained substantial interest in the computer vision community Han et al. (2020; 2021); Fini et al. (2021); Wen et al. (2023); Jia et al. (2021); Zhao & Han (2021). The task is to leverage knowledge from a number of labelled images, in order to discover and cluster images from novel classes in unlabelled data. Such a task naturally occurs in many practical settings; from products in a supermarket, to animals in the wild, to street objects for an autonomous vehicle. Specifically, Generalized Category Discovery (GCD) Vaze et al. (2022) has recently emerged as a challenging variant of the problem in which the unlabelled data can contain both instances from ‘seen’ and ‘unseen’ classes. As such, the problem is succinctly phrased as: “given a dataset, some of which is labelled, categorise all unlabelled instances (whether or not they come from labelled classes)”.
In this paper, we challenge a key, but often ignored, assumption in this setting: GCD methods still assume that all instances in the unlabelled set come from the same domain as the labelled data. In practise, unlabelled images may not only contain novel categories, but also exhibit low-level covariate shift Sun et al. (2022); Yan et al. (2019). It has long been established that the performance of image classifiers degrades substantially in the presence of such shifts Ganin et al. (2016); Tzeng et al. (2014); Zhang et al. (2019) and, indeed, we find that existing GCD models perform poorly in such a setting. Compared to related literature in, for instance, domain adaptation Du et al. (2021); Chen et al. (2022b); Zhu et al. (2023) or domain generalization Shi et al. (2022); Harary et al. (2022) the task proposed here presents a dual challenge: models must be robust to the low-level covariate shift while remaining sensitive to semantic novelty.
@@
To tackle these problems, we introduce the ‘HiLo’ architecture and learning framework. The HiLo architecture extracts both ‘low-level’ (early layer) and ‘high-level’ (late layer) features from a vision transformer Dosovitskiy et al. (2020). While extracting features at multiple stages of the network has been performed in domain adaptation Bousmalis et al. (2016); Peng et al. (2019b); Liu et al. (2020), we further introduce an explicit loss term to minimise mutual information between the two sets of features (Section 3.2.1).
@@
The intuition is that the covariate and semantic information in the data is (by definition) independent, and that the inductive bias of deep architectures is likely to represent low-level covariate information in early layers, and abstract semantic information in later ones Olah et al. (2017); Zhou et al. (2021). Next, we take inspiration from a strong method from the domain adaptation field, PatchMix Zhu et al. (2023), which works by performing mixup augmentation in the embedding space of a pretrained transformer. While naive application of this method does not account for semantic novelty in unlabelled data, we extend the PatchMix objective to allow training with both a self-supervised contrastive objective (Section 3.2.2), and a semantic clustering loss (Section 3.2.2). With these changes, the PatchMix style augmentation is tailored to leverage both the labelled and unlabelled data available in the GCD setting. Our ‘HiLo’ feature design in our framework enables the model to disentangle domain and semantic features, while patch mixing allows the model to bridge the domain gap among images and focus more on determining the semantic shifts. Therefore, we introduce the patch mixing idea into our ‘HiLo’ framework, equipping it with a strong capability to discover novel categories from unlabelled images in the presence of domain shifts.
@@
Finally, we find that curriculum learning Bengio et al. (2015); Zhou et al. (2020); Wu & Vorobeychik (2022) is particularly applicable to the setting introduced in this work (Section 3.2.3). Specifically, the quality of the learning signal differs substantially across different partitions of the data: from a clean supervised signal on the labelled set; to unsupervised signals from unlabelled data which may or may not come from the same domain and categories. It is non-trivial to train a GCD model to discover novel categories in the presence of both domain shifts and semantic shifts in the unlabelled data. To address this challenge, we introduce a curriculum learning approach which gradually increases the sampling probability weight of samples predicted as from unknown domains, as training proceeds. Our sequential learning process prioritizes the discovery of semantic categories initially and progressively enhances the model’s ability to handle covariate shifts, which cannot be achieved by simply adopting existing domain adaptation methods.
@@
To evaluate out models, we construct the ‘SSB-C’ benchmark suite – based on the recent Semantic Shift Benchmark (SSB) Vaze et al. (2021) – with domain shifts introduced by synthetic corruptions following ImageNet-C Hendrycks & Dietterich (2019). On this benchmark, as well as on a large-scale DomainNet evaluation with real data Peng et al. (2019a), we also reimplement a range of performant baselines from the category discovery literature. We find that, on both benchmarks, our method substantially outperforms all existing category discovery models Vaze et al. (2022); Wen et al. (2023); Han et al. (2019); Fini et al. (2021).
@@
In summary, we make the following key contributions: (i) We formalize a challenging open-world task for category discovery in the presence of domain shifts; (ii) We develop a new method, HiLo, which disentangles covariate and semantic features to tackle the problem, extending state-of-the-art methods from the domain adaptation literature; (iii) We reimplement a range of category discovery models on a benchmark suite containing both fine-grained and coarse-grained datasets, with real and synthetic corruptions. (iv) We demonstrate that, on all datasets, our method substantially outperforms current state-of-the-art category discovery methods with finetuned hyperparameters.
@@ 2 Related work
Category discovery was firstly studied as novel category discovery (NCD) Han et al. (2019) and recently extended to generalized category discovery (GCD) Vaze et al. (2022). GCD extends NCD by including unlabelled images from both labelled and novel categories. Many successful NCD methods have been proposed (e.g., DTC Han et al. (2019), RankStats Han et al. (2020; 2021), WTA Jia et al. (2021), DualRank Zhao & Han (2021), OpenMix Zhong et al. (2021b), NCL Zhong et al. (2021a), UNO Fini et al. (2021), class-relation knowledge distillation Gu et al. (2023)), they do not address domain shifts. Recent work Zang et al. (2023) considers domain shifts in NCD with labelled target domain images. We focus on GCD without any labelled instances from new domains, where unlabelled images may come from multiple novel domains. There is a growing body of literature for category discovery, including non-parametric methods (e.g., GCD Vaze et al. (2022), CiPR Hao et al. (2023)); parametric methods (e.g., SimGCD Wen et al. (2023)), ORCA Cao et al. (2021), $\mu$-GCD Vaze et al. (2023); prompt based techniques (e.g., PromptCAL Zhang et al. (2023), SPTNet Wang et al. (2024a), PromptGCD Cendra et al. (2024)), debiased learning (e.g., DebGCD Liu & Han (2025)), and other specialized methods (e.g., GPC Zhao et al. (2023), DCCL Pu et al. (2023), InfoSieve Rastegar et al. (2024)). However, existing GCD methods do not consider domain shifts in unlabelled data.
Semi-supervised learning (SSL) aims to develop robust classification models using both labelled and unlabelled data, assuming instances belong to the same class set. Consistency-based approaches, such as Mean-teacher Tarvainen & Valpola (2017), Mixmatch Berthelot et al. (2019), and Fixmatch Sohn et al. (2020), have demonstrated effectiveness in SSL. Recent methods Chen et al. (2020b; c; 2021) have enhanced SSL performance by incorporating contrastive learning (e.g., Chen et al. (2020a)He et al. (2020)). Several studies Wang et al. (2022); Rizve et al. (2022); Wang et al. (2024b); Sun et al. (2024) have extended standard SSL to open-world settings.
@@
Unsupervised domain adaptation (UDA) adapts models from a source domain to a target domain, with labelled data from the former and unlabelled data from the latter. UDA methods are categorized into moment matching Tzeng et al. (2014); Long et al. (2015; 2017); Zhang et al. (2019) and adversarial learning Ganin et al. (2016); Gao et al. (2021); Tang & Jia (2020) methods. DANN, FGDA, DADA are popular examples using a min-max game. MCD and SWD implicitly use adversarial learning with $L_{1}$ distance and sliced Wasserstein discrepancy, respectively. CGDM Du et al. (2021) leverages cross-domain gradient discrepancy, while Chen et al. (2022b) couples NWD with a single task-specific classifier with implicit K-Lipschitz constraint. PMTrans Zhu et al. (2023) aligns the source and target domains with the intermediate domain by employing semi-supervised mixup losses in both feature and label spaces. MCC Jin et al. (2020) minimizes between-class confusion and maximizing within-class confusion, while NWD Chen et al. (2022b) uses a single task-specific classifier with implicit K-Lipschitz constraint to obtain better robustness for all the domain adaptation scenarios.
@@ 3 HiLo networks for GCD with domain shifts
Problem statement. We define Generalized Category Discovery with domain shifts as the task of classifying images from mixed domains $\Omega=\Omega^{a}\cup\Omega^{b}$ (where $\Omega^{a}\cap\Omega^{b}=\emptyset$ and $\Omega^{b}$ may contain multiple domains in practise), only having access to partially labelled samples from domain $\Omega^{a}$. The goal is to assign class labels to the remaining images, whose categories and domains may be seen or unseen in the labelled images. Formally, let $\mathcal{D}$ be an open-world dataset consisting of a labelled set $\mathcal{D}^{l}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N_{l}}\subset\mathcal{X}^{l}\times\mathcal{Y}^{l}$ and an unlabelled set $\mathcal{D}^{u}=\{\bm{x}_{i}\}_{i=1}^{N_{u}}\subset\mathcal{X}^{u}$. The label space for labelled samples is $\mathcal{Y}^{l}=\mathcal{C}_{1}$ and for unlabelled samples is $\mathcal{Y}^{u}=\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$, where $\mathcal{C}$, $\mathcal{C}_{1}$, and $\mathcal{C}_{2}$ represent the label sets for ‘All’, ‘Old’, and ‘New’ categories, respectively. It is important to note that $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$. The objective of GCD with domain shifts is to classify all unlabelled images in $\mathcal{D}^{u}$ (from either $\Omega^{a}$ or $\Omega^{b}$) using only the labels in $\mathcal{D}^{l}$. This is different from the setting of NCD with domain shift Yu et al. (2022), which assumes $\mathcal{Y}^{l}\cap\mathcal{Y}^{u}=\emptyset$, and GCD, which assumes $\Omega^{a}=\Omega^{b}$ with $|\Omega^{a}|=|\Omega^{b}|=1$. For notation simplicity, hereafter we omit the subscript $i$ for each image $\bm{x}_{i}$.
@@ 3.1 Background: SimGCD
SimGCD Wen et al. (2023) is a representative end-to-end baseline for GCD, which integrates two primary losses for representation learning and parametric classification: (1) a contrastive loss $\mathcal{L}^{rep}$ based on InfoNCE Oord et al. (2018) is applied for the representation learning of the feature backbone; and (2) a cross-entropy loss $\mathcal{L}^{cls}$ for training a cosine classification head Gidaris & Komodakis (2018), utilizing different image views as pseudo-labels for one another. Following Vaze et al. (2022), SimGCD employs the ViT model as the backbone containing $m$ Transformer layers. Let $\mathcal{F}$ be the feature extractor consisting of these $m$ layers and $\mathcal{H}$ be a projection head. For an input image $\bm{x}$, a $\ell_{2}$-normalised feature can be obtained by $\bm{z}=\mathcal{H}(\mathcal{F}(\varphi(\bm{x})))$, where $\varphi$ is a standard embedding layer before the multi-head attention layers in the ViT model. The representation loss is
$\mathcal{L}^{rep}(\bm{x})=-\frac{1}{|\mathcal{P}(\bm{x})|}\sum_{\bm{z}^{+}\in\mathcal{P}(\bm{x})}\log\sigma(\bm{z}\cdot\bm{z}^{+};\tau),$ (1)
@@
where $\sigma(\cdot;\tau)$ is the softmax operation with a temperature $\tau$ for scaling and $\mathcal{P}(\bm{x})$ denotes the positive feature set for each $\bm{x}$. Suppose we sample a batch $\mathcal{B}$, which contains labelled images and unlabelled images, denoted as $\mathcal{B}^{l}$ and $\mathcal{B}^{u}$, respectively. For each $\bm{x}\in\mathcal{B}$ (either a labelled or unlabelled image), $\mathcal{P}(\bm{x})$ contains only the feature of a different view of the same image. For each $\bm{x}\in\mathcal{B}^{l}$, an additional $\mathcal{P}(\bm{x})$ including features of other images from the same class and the feature of a different view of the same image is also used for supervised contrastive learning. Likewise, the classification loss can be written as
$\mathcal{L}^{cls}(\bm{x})=-\sum_{\bm{w}\in\bm{W}}\bm{q}\log\sigma(\hat{\bm{z}}\cdot\bm{w};\tau),$ (2)
@@
where $\bm{W}$ is a set of prototypes and each vector $\bm{w}$ in $\bm{W}$ represents a $\ell_{2}$-normalised learnable class prototype. $\hat{\bm{z}}$ is the $\ell_{2}$-normalised vector of $\mathcal{F}(\varphi(\bm{x}))$. For each $\bm{x}\in\mathcal{B}$, $\bm{q}$ is the pseudo-label from a sharpened prediction of a different view of the same image. For each $\bm{x}\in\mathcal{B}^{l}$, an additional $\bm{q}$ as the one-hot ground-truth vector is also used for supervised learning. Let $\mathcal{L}^{r,c}$ be the summation of $\mathcal{L}^{rep}$ and $\mathcal{L}^{cls}$ for simplification, the overall loss can then be written as:
$\mathcal{L}_{sim}=\lambda\sum_{\bm{x}\in\mathcal{B}}\mathcal{L}^{r,c}(\bm{x})+(1-\lambda)\sum_{\bm{x}\in\mathcal{B}^{l}}\mathcal{L}^{r,c}(\bm{x})+\epsilon\Delta,$ (3)
@@
where $\mathcal{B}^{l}$ denotes the subset of labelled samples in the current mini-batch, and $\Delta$ is an entropy maximization term to prevent pseudo-label collapse Assran et al. (2022). Finally, $\lambda$ and $\epsilon$ are hyperparameters, and we refer to the original work for further details Wen et al. (2023).